OpenStudy (anonymous):

For the titration of 25.0mL of 0.20M hydrofluoric acid with 0.20M NaOH, determine the volume of base added when pH is (a)2.85 (b)3.15 (c)11.89?

OpenStudy (anonymous):

can i tell you for the first part only? and you can solve the same way for the next parts?

OpenStudy (anonymous):

tell me how i can approach all 3

OpenStudy (anonymous):

ok.. dude! wait!

OpenStudy (anonymous):

first part : pH = 2.85 i.e. -log(H+) = 2.85. so, [H+] = 1.4*10^-3 this is the resultant [H+] ion concentration.. by mixing 0.025L of HF and say V litres of NaOH. so. the resultant number of moles of H+ = concentration * volume = 1.4*10^-3 *(0.025 +V) and, HF dissociates as:- \[HF \leftarrow \rightarrow H ^{+}+F ^{-}\] 0.2 0 0 ...............(initially) 0.2-X x x ..................(at time t) and we know, \[K _{a}(HF) = \frac{ x ^{2} }{ 0.2-x }\] Ka(HF) = 3.47 *10^-4 solving this for x, we get x = 0.00816. thus, [H+] = x = 8.16*10^-3 NUMBER OF MOLES OF H+ FROM ACID = CONC. * VOLUME 8.16*10^-3 * 0.025 = 2.04*10^-4 number of moles of OH- = conc. of base * volume taken = 0.2*V as this no. of moles of OH- react with H+ , we will subtract this 0.2V from 2.04*10^-4 and put it equal to 1.4*10^-3(0.025+V) solving this for V, you will get the answer. i.e. you need to solve the equation. 2.04*10^-4 - 0.2V = 1.4*10^-3 ( 0.025 + V) the ans. will come out to be 8.3 *10^-4 Litres.. i think

OpenStudy (anonymous):

see if this helps.

OpenStudy (anonymous):

for first part ans = 0.837 mL for seconddd part ans == 0.928 mL third padrt 1.02 ml are these the answers?????????????????

OpenStudy (anonymous):

are these the answers?